# Can GitHub Issues Help in App Review Classifications?

### v2

@@ 3.2.2 Extracting Target Information
For example, Figure 1a illustrates a bug issue with a specific template comprising several main sections, namely the First Occurred, Steps to Reproduce, Expected Behaviour, and Actual Behaviour. The first section contains the date of occurrence of the bug, aiding developers in locating it. The second section lists the detailed steps for reproducing the bug. Typically, users only state an overview of encountered bugs and refer to the last steps in most reviews [36]. As we see in Figure 1b, a user has mentioned the last three steps without providing many details for replicating the bug. The third section also specifies the desired behavior of the user, which we found to often have the same content as the Actual Behaviour section, albeit using a positive verb. This section can lead to errors in the final model. So, the content of the first three sections in the issue shown in Figure 1a provides the details for developers to reproduce and fix this bug, while user reviews do not include this type of information. Therefore, these sections cannot be helpful for us, and we remove them. On the other hand, the content of the Actual Behaviour section is suitable for our purpose. It describes the current behavior of the app and provides an overview of the bug, similar to a user review (Figure 1b). As a result, we need to choose sections that provide a general description of an issue, similar to user reviews, to create an auxiliary dataset. During this phase, we manually analyze the issue bodies and templates to identify the necessary information. Subsequently, we define 19 linguistic patterns that allow us to extract this relevant information accurately and efficiently.
Developers have the option to either follow a specific template for submitting issues or write them freely and informally. Issue templates usually contain certain sections that provide valuable information for a particular purpose. These templates aid in creating structured content and assist issue authors in recalling the necessary information [37]. Issue templates are stored in the “.github/ISSUE_TEMPLATE” folder as “.md”, “.yaml”, or “.yml” files [38]. These files contain instructions for preparing appropriate content for different sections of a specific issue. By utilizing these templates, we can identify sections that offer target information compatible with user reviews. In our research, we collected a total of 645 available issue templates from 291 repositories using the GitHub REST API.
@@
Figure 6: The Number of Issue Templates in Five Groups: Bug, Feature, Other, Issue, and Deleted
(a) A Bug Template in Md format from the AdAway Repository [39]
(b) A Bug Template in YAML format from the VocableTrainer-Android Repository [40]
Figure 7: Two Samples of Bug Template in Yaml and Md Format.
@@
Table 5: Samples of the Target Information Extracted from Issue Bodies
@@
Table 6: Linguistic Patterns (constituent words represented in lemmatized format. B: Bug Report, F: Feature Request, O: Other)
@@
P5 r“problem (?:statement\$|address\$)|(?:bug|tell us about) (?:problem\$)”
@@
P9 r“.*featur (?:request|suggest\$)|.*suggest (?:featur\$)”
@@
P11 r“what (?:issu\$|about\$)|what issu about”
@@
Table 7: Examples of Linguistic Patterns
@@
3.2.3 Preprocessing Textual Data
@@
  1. 1.
    Filtering textual data: The body of issues is usually processed by removing additional information that has less value for our purpose, which increases the accuracy of the proposed approach [15]. We remove noisy contents such as checklists, tags, links, code fragments, variable names, error messages, mention terms, and issue numbers. Additionally, we exclude special phrases listed in Table 8, as well as any phrases that start and end with an underline symbol (such as “_created by X_”). These phrases do not provide helpful information and usually refer to the author or source of the issue.
  2. 2.
@@
  5. 5.
    Transforming negative modifiers to a single form “not”: Negative terms play a crucial role in determining the intention of a review [34]. So, we identify 44 negative modifiers (Table 8) and replace them with “not” to reduce the number of features.
  6. 6.
    Removing stop words: we use the NLTK stop words, with exceptions for negative modifiers listed in Table 8 and three specific words “Could”, “Would”, and “Should”. The exclusion of the former is explained in the previous step. Also, the latter represent important keywords for indicating the review intention [41]. Additionally, the term “have to” (similar to “must”) can be effective in predicting a feature request review. As this term consists of two words, “have” and “to,” they are identified as stop words. Consequently, we convert it to the term “have-to” to prevent its deletion during preprocessing.
  7. 7.
@@
  8. 8.
    Removing non-informative items: After applying the above steps to the titles and bodies, we remove issues that contain less than three words because they provide little information [42]. Additionally, we filter out issues with variable names in the title because our investigation shows they are often related to implementation topics.
@@
Table 8: Special Words Utilized in Preprocessing Steps
@@
3.3 Preparing the Primary Dataset
As mentioned above, previous studies classified user reviews from various aspects, with the most frequent being selected intentions such as bug reports and feature requests. They often labeled a dataset manually through a standard process to train and test their proposed model. In this paper, we utilize five well-known labeled review datasets provided by Guzman et al. [6], Maalej et al. [1], Jha et al. [33], Gu et al. [31], and Scalabrino et al. [34] as primary datasets. It is worth noting that the Jha et al. dataset contains some reviews from Maalej et al. We removed these duplicates from the Jha et al. dataset to avoid redundancy, as we use both datasets separately.
There are several reasons for choosing these five datasets. First, they are publicly available and have been labeled through a suitable process by human experts. Second, these datasets have been used in similar research studies [7, 22], making them well-established and reliable sources for comparison and evaluation. Therefore, studying their generalizability improvements through augmentation with the help of an auxiliary dataset from another source can be highly beneficial. Third, the review classes in these datasets can be adapted to our intentions. Since these datasets use different taxonomies or label names, we unified related classes to the bug report, feature request, and other intentions according to their definitions. Additionally, some classes were inconsistent with our intentions or could not be assigned to just one class, so we excluded them. For example, reviews related to the security label in the Scalarbino et al. dataset could include bug reports or requests for improving security-related app features. We present the adapted classes in Table 9. Forth, selected datasets were collected from various apps with different data sizes. This way, we can analyze the feasibility of augmenting review datasets with an auxiliary dataset derived from GitHub issues under different conditions.
@@ 3.3 Preparing the Primary Dataset
Table 9: Adapted Classes in Primary Datasets
@@
PD1 Guzman [6] Bug report User request Praise, Usage scenario Feature shortcoming, Feature strength, Complaint, Noise
PD2 Maalej [1] Bug Feature UserExperience, Rating -
PD3 Jha [33] BugReport FeatureRequest Other -
PD4 Gu [31] bug reports feature requests praise, others aspect evaluation
PD5 Scalabrino [34] BUG FEATURE OTHER PERFORMANCE, SECURITY, ENERGY, USABILITY
Table 10: Details of Primary Datasets Used in Our Approach
@@
3.4 Dataset Augmentation
As mentioned earlier, our goal is to enhance the performance of review classification models by augmenting the training dataset with labeled data extracted from GitHub issues. To this end, we prepared an auxiliary dataset of GitHub issues by applying various preprocessing steps. Also, we utilized five manually labeled review datasets as primary datasets and adapted their labels to align with our intentions (bug report, feature request, other). During this phase, we integrate these two types of datasets based on the three methods of Between-Apps, Within-Apps, and Within-Context Analysis. Maalej et al. [1] introduced the first two, and we incorporate these methods in our approach based on their definition.
@@ 3.4 Dataset Augmentation
Another method is Within-Apps Analysis, which has received less attention because of the labeling challenges it presents. Based on this method, review classifiers are trained for each app using a specifically labeled dataset. In the Within-Apps Analysis, we use the processed issues of the same app to prepare an auxiliary dataset. This method allows us to leverage app-specific information and characteristics, tailoring the classifier to each app’s unique review requirements. Indeed, obtaining a sufficient amount of labeled data within each app presents a significant challenge that needs to be addressed to ensure the effectiveness of the Within-Apps Analysis approach. This is particularly true for many Android apps, especially those that are not popular and may have limited labeled issues available. Hassan et al. [43] demonstrated that analyzing apps while considering the data provided by similar apps can lead to helpful information and reveal other aspects. Therefore, we propose the Within-Context Analysis as an alternative method to overcome the limitation of app-specific labelled datasets. This method leverages the context data of an app to prepare its auxiliary dataset. Instead of relying solely on app-specific labeled issues, we utilize GitHub issues from similar apps with similar characteristics and functionalities. By doing so, we can gather the required information for the target app from a broader range of sources, enhancing the potential for a more comprehensive and contextually relevant review classification. Additionally, we use the repository descriptions in GitHub to find similarities among apps, as detailed in Section 3.4.1.
@@
3.4.1 Finding Similar Repositories
Figure 8: An Example of a GitHub Repository
In this step, we create a profile for each repository to find similar ones [44]. We utilize the information available in the ReadMe file and the About section of repositories (as shown in Figure 8). The ReadMe file contains descriptions of the repository, including its purpose, features, usage instructions, contribution guidelines, etc. We manually analyzed 200 randomly selected ReadMe files to extract the necessary information. Our findings revealed that the initial section of the ReadMe file and sections titled introduction, description, features, what it does, about, about the project, overview, summary, todo often describe the repository’s purpose and features. Furthermore, the About section provides a concise app description. We extract the content of these sections to construct the app profiles and preprocess them using steps 2 to 8 in Section 3.2.3. We employ tf-idf [45] to transform each app profile into a vector and calculate cosine similarities to find similar apps.
3.5 Review Classification
As previously mentioned, we utilize processed issues, including issue titles and content of specific sections within issue bodies, to augment manually labeled reviews. These are integrated to generate augmented datasets, resulting in more accurate predictions of user review intentions. Recent studies have demonstrated transformer-based models performed well in classifying user reviews [22, 46, 47]. In this paper, we employ our augmented datasets to fine-tune four pre-trained models, BERT, RoBERTa, DistilBERT, and ALBERT.
We aim to classify user reviews into three intention classes: bug reports, feature requests, and others. Consequently, we face a multi-class classification problem with two alternatives (i) training a binary classifier for each class or (ii) training a multi-class classifier [1]. Since a review could fall simultaneously into the bug report and feature request categories, we choose the former in training classifiers. Also, prior studies [1, 22] found that training binary classifiers provide better accuracy than multi-class classifiers. On the other hand, considering that the others class comprises reviews that do not pertain to bug reports or feature requests, it becomes apparent that training a distinct classifier for this category is unnecessary. As a result, we train two classifiers specifically for predicting the bug report and feature request classes.
4 Experiment Settings
4.1 Research Questions
@@ 4.2 Dataset and Models
To create a truth dataset, we followed the labeling process steps used by Guzman et al. [6]. First, we prepared a guideline containing the definitions of each class along with several examples. This guideline aided the annotators in comprehending the process better and minimizing disagreements. Second, we randomly selected 200 reviews from each app, including reviews from the bug report, feature request, and other classes. Four students, with an average programming experience of 9.5 years, participated as annotators in the labeling process. We assigned the selected reviews to them so that two annotators assessed each review. They manually analyzed these reviews and selected proper labels for them. Overall, they agreed on 91% of the labels and achieved a 0.79 Cohen’s kappa [48], providing reliable results. Finally, we investigated disputed reviews and held a meeting to make final decisions about them. Annotators discussed each other during this meeting and agreed on the final labels. Table 11 shows the specifications of the labeled truth dataset.
Table 11: Details of Truth Datasets (Revs=Reviews)
Apps #Total Revs #Bug Revs #Feature Revs #Other Revs #Total Issues #Bug Issues #Feature Issues #Other Issues ownCloud [49, 50] 200 90 63 55 3,086 2,560 527 293 AntennaPod [24, 23] 200 33 58 109 160 45 107 55 MetaMask [51, 52] 200 115 49 40 159 91 70 30 Firefox Nightly [53, 54] 200 38 84 82 6,077 5,243 854 346 Firefox Focus [55, 56] 200 37 78 87 735 597 139 44 SUM 1000 313 332 373 10,217 8,536 1,697 768
We employed the TFAutoModelForSequenceClassification class [57] from Huggingface’s Transformers library to fine-tune pre-trained models. We trained the classification models for 3 epochs using a batch size of 16 and a learning rate of 5e-5 for the Adam optimizer [58]. Furthermore, we set the maximum input string length to 50, consistent with our datasets. All models were executed on Google Colab, which offers a free GPU with 12.7 GB of system RAM, 15.0 GB of GPU RAM, and 78.2 GB of disk storage. Also, our investigation, which corresponds to RQ4, demonstrates that the DistilBERT model achieves the highest F1-score in the average of the bug report and feature request results with the shortest time for training and prediction phases. So, we used it to train different review classifiers in preceding experiments.
4.3 Evaluation Measures

### v3

@@ 3.2.2. Extracting Target Information
For example, Figure 1(a) illustrates a bug issue with a specific template comprising several main sections, namely the First Occurred, Steps to Reproduce, Expected Behaviour, and Actual Behaviour. The first section contains the date of occurrence of the bug, aiding developers in locating it. The second section lists the detailed steps for reproducing the bug. Typically, users only state an overview of encountered bugs and refer to the last steps in most reviews (Borges et al., 2020). As we see in Figure 1(b), a user has mentioned the last three steps without providing many details for replicating the bug. The third section also specifies the desired behavior of the user, which we found to often have the same content as the Actual Behaviour section, albeit using a positive verb. This section can lead to errors in the final model. So, the content of the first three sections in the issue shown in Figure 1(a) provides the details for developers to reproduce and fix this bug, while user reviews do not include this type of information. Therefore, these sections cannot be helpful for us, and we remove them. On the other hand, the content of the Actual Behaviour section is suitable for our purpose. It describes the current behavior of the app and provides an overview of the bug, similar to a user review (Figure 1(b)). As a result, we need to choose sections that provide a general description of an issue, similar to user reviews, to create an auxiliary dataset. During this phase, we manually analyze the issue bodies and templates to identify the necessary information. Subsequently, we define 19 linguistic patterns that allow us to extract this relevant information accurately and efficiently.
Developers have the option to either follow a specific template for submitting issues or write them freely and informally. Issue templates usually contain certain sections that provide valuable information for a particular purpose. These templates aid in creating structured content and assist issue authors in recalling the necessary information (Li et al., 2023). Issue templates are stored in the “.github/ISSUE_TEMPLATE” folder as “.md”, “.yaml”, or “.yml” files (Nikeghbal et al., 2024). These files contain instructions for preparing appropriate content for different sections of a specific issue. By utilizing these templates, we can identify sections that offer target information compatible with user reviews. In our research, we collected a total of 645 available issue templates from 291 repositories using the GitHub REST API.
@@
Figure 6. The Number of Issue Templates in Five Groups: Bug, Feature, Other, Issue, and Deleted
Figure 7. Two Samples of Bug Template in Yaml and Md Format.
@@
Table 5. Samples of the Target Information Extracted from Issue Bodies
@@
Table 6. Linguistic Patterns (constituent words represented in lemmatized format. B: Bug Report, F: Feature Request, O: Other)
@@
P5 r“problem (?:statement\$|address\$)—(?:bug|tell us about) (?:problem\$)”
@@
P9 r“.*featur (?:request|suggest\$)—.*suggest (?:featur\$)”
@@
P11 r“what (?:issu\$—about\$)|what issu about”
@@
Table 7. Examples of Linguistic Patterns
@@
3.2.3. Preprocessing Textual Data
@@
  1. (1)
    Filtering textual data: The body of issues is usually processed by removing additional information that has less value for our purpose, which increases the accuracy of the proposed approach (Izadi et al., 2022). We remove noisy contents such as checklists, tags, links, code fragments, variable names, error messages, mention terms, and issue numbers. Additionally, we exclude special phrases listed in Table 8, as well as any phrases that start and end with an underline symbol (such as “_created by X_”). These phrases do not provide helpful information and usually refer to the author or source of the issue.
  2. (2)
@@
  5. (5)
    Transforming negative modifiers to a single form “not”: Negative terms play a crucial role in determining the intention of a review (Scalabrino et al., 2019). So, we identify 44 negative modifiers (Table 8) and replace them with “not” to reduce the number of features.
  6. (6)
    Removing stop words: we use the NLTK stop words, with exceptions for negative modifiers listed in Table 8 and three specific words “Could”, “Would”, and “Should”. The exclusion of the former is explained in the previous step. Also, the latter represent important keywords for indicating the review intention (Maalej and Nabil, 2015). Additionally, the term “have to” (similar to “must”) can be effective in predicting a feature request review. As this term consists of two words, “have” and “to,” they are identified as stop words. Consequently, we convert it to the term “have-to” to prevent its deletion during preprocessing.
  7. (7)
@@
  8. (8)
    Removing non-informative items: After applying the above steps to the titles and bodies, we remove issues that contain less than three words because they provide little information (Assi et al., 2021). Additionally, we filter out issues with variable names in the title because our investigation shows they are often related to implementation topics.
@@
Table 8. Special Words Utilized in Preprocessing Steps
@@
3.3. Preparing the Primary Dataset
As mentioned above, previous studies classified user reviews from various aspects, with the most frequent being selected intentions such as bug reports and feature requests. They often labeled a dataset manually through a standard process to train and test their proposed model. In this paper, we utilize five well-known labeled review datasets provided by Guzman et al. (Guzman et al., 2015), Maalej et al. (Maalej et al., 2016), Jha et al. (Jha and Mahmoud, 2018), Gu et al. (Gu and Kim, 2015), and Scalabrino et al. (Scalabrino et al., 2019) as primary datasets. It is worth noting that the Jha et al. dataset contains some reviews from Maalej et al. We removed these duplicates from the Jha et al. dataset to avoid redundancy, as we use both datasets separately.
There are several reasons for choosing these five datasets. First, they are publicly available and have been labeled through a suitable process by human experts. Second, these datasets have been used in similar research studies (Devine et al., 2023; Hadi and Fard, 2021), making them well-established and reliable sources for comparison and evaluation. Therefore, studying their generalizability improvements through augmentation with the help of an auxiliary dataset from another source can be highly beneficial. Third, the review classes in these datasets can be adapted to our intentions. Since these datasets use different taxonomies or label names, we unified related classes to the bug report, feature request, and other intentions according to their definitions. Additionally, some classes were inconsistent with our intentions or could not be assigned to just one class, so we excluded them. For example, reviews related to the security label in the Scalarbino et al. dataset could include bug reports or requests for improving security-related app features. We present the adapted classes in Table 9. Forth, selected datasets were collected from various apps with different data sizes. This way, we can analyze the feasibility of augmenting review datasets with an auxiliary dataset derived from GitHub issues under different conditions.
@@ 3.3. Preparing the Primary Dataset
Table 9. Adapted Classes in Primary Datasets
@@
PD1 Guzman (Guzman et al., 2015) Bug report User request Praise, Usage scenario Feature shortcoming, Feature strength, Complaint, Noise
PD2 Maalej (Maalej et al., 2016) Bug Feature UserExperience, Rating -
PD3 Jha (Jha and Mahmoud, 2018) BugReport FeatureRequest Other -
PD4 Gu (Gu and Kim, 2015) bug reports feature requests praise, others aspect evaluation
PD5 Scalabrino (Scalabrino et al., 2019) BUG FEATURE OTHER PERFORMANCE, SECURITY, ENERGY, USABILITY
Table 10. Details of Primary Datasets Used in Our Approach
@@
3.4. Dataset Augmentation
As mentioned earlier, our goal is to enhance the performance of review classification models by augmenting the training dataset with labeled data extracted from GitHub issues. To this end, we prepared an auxiliary dataset of GitHub issues by applying various preprocessing steps. Also, we utilized five manually labeled review datasets as primary datasets and adapted their labels to align with our intentions (bug report, feature request, other). During this phase, we integrate these two types of datasets based on the three methods of Between-Apps, Within-Apps, and Within-Context Analysis. Maalej et al. (Maalej et al., 2016) introduced the first two, and we incorporate these methods in our approach based on their definition.
@@ 3.4. Dataset Augmentation
Another method is Within-Apps Analysis, which has received less attention because of the labeling challenges it presents. Based on this method, review classifiers are trained for each app using a specifically labeled dataset. In the Within-Apps Analysis, we use the processed issues of the same app to prepare an auxiliary dataset. This method allows us to leverage app-specific information and characteristics, tailoring the classifier to each app’s unique review requirements. Indeed, obtaining a sufficient amount of labeled data within each app presents a significant challenge that needs to be addressed to ensure the effectiveness of the Within-Apps Analysis approach. This is particularly true for many Android apps, especially those that are not popular and may have limited labeled issues available. Hassan et al. (Hassan et al., 2022) demonstrated that analyzing apps while considering the data provided by similar apps can lead to helpful information and reveal other aspects. Therefore, we propose the Within-Context Analysis as an alternative method to overcome the limitation of app-specific labelled datasets. This method leverages the context data of an app to prepare its auxiliary dataset. Instead of relying solely on app-specific labeled issues, we utilize GitHub issues from similar apps with similar characteristics and functionalities. By doing so, we can gather the required information for the target app from a broader range of sources, enhancing the potential for a more comprehensive and contextually relevant review classification. Additionally, we use the repository descriptions in GitHub to find similarities among apps, as detailed in Section 3.4.1.
@@
3.4.1. Finding Similar Repositories
Figure 8. An Example of a GitHub Repository
In this step, we create a profile for each repository to find similar ones (Li et al., 2017). We utilize the information available in the ReadMe file and the About section of repositories (as shown in Figure 8). The ReadMe file contains descriptions of the repository, including its purpose, features, usage instructions, contribution guidelines, etc. We manually analyzed 200 randomly selected ReadMe files to extract the necessary information. Our findings revealed that the initial section of the ReadMe file and sections titled introduction, description, features, what it does, about, about the project, overview, summary, todo often describe the repository’s purpose and features. Furthermore, the About section provides a concise app description. We extract the content of these sections to construct the app profiles and preprocess them using steps 2 to 8 in Section 3.2.3. We employ tf-idf (Salton and Buckley, 1988) to transform each app profile into a vector and calculate cosine similarities to find similar apps.
3.5. Review Classification
As previously mentioned, we utilize processed issues, including issue titles and content of specific sections within issue bodies, to augment manually labeled reviews. These are integrated to generate augmented datasets, resulting in more accurate predictions of user review intentions. Recent studies have demonstrated transformer-based models performed well in classifying user reviews (Hadi and Fard, 2021; Henao et al., 2021; Araujo et al., 2022). In this paper, we employ our augmented datasets to fine-tune four pre-trained models, BERT, RoBERTa, DistilBERT, and ALBERT.
We aim to classify user reviews into three intention classes: bug reports, feature requests, and others. Consequently, we face a multi-class classification problem with two alternatives (i) training a binary classifier for each class or (ii) training a multi-class classifier (Maalej et al., 2016). Since a review could fall simultaneously into the bug report and feature request categories, we choose the former in training classifiers. Also, prior studies (Maalej et al., 2016; Hadi and Fard, 2021) found that training binary classifiers provide better accuracy than multi-class classifiers. On the other hand, considering that the others class comprises reviews that do not pertain to bug reports or feature requests, it becomes apparent that training a distinct classifier for this category is unnecessary. As a result, we train two classifiers specifically for predicting the bug report and feature request classes.
4. Experiment Settings
4.1. Research Questions
@@ 4.2. Dataset and Models
To create a truth dataset, we followed the labeling process steps used by Guzman et al. (Guzman et al., 2015). First, we prepared a guideline containing the definitions of each class along with several examples. This guideline aided the annotators in comprehending the process better and minimizing disagreements. Second, we randomly selected 200 reviews from each app, including reviews from the bug report, feature request, and other classes. Four students, with an average programming experience of 9.5 years, participated as annotators in the labeling process. We assigned the selected reviews to them so that two annotators assessed each review. They manually analyzed these reviews and selected proper labels for them. Overall, they agreed on 91% of the labels and achieved a 0.79 Cohen’s kappa (Cohen, 1960), providing reliable results. Finally, we investigated disputed reviews and held a meeting to make final decisions about them. Annotators discussed each other during this meeting and agreed on the final labels. Table 11 shows the specifications of the labeled truth dataset.
Table 11. Details of Truth Datasets (Revs=Reviews)
Apps #Total Revs #Bug Revs #Feature Revs #Other Revs #Total Issues #Bug Issues #Feature Issues #Other Issues ownCloud (own, 2023b, a) 200 90 63 55 3,086 2,560 527 293 AntennaPod (Ate, 2023b, a) 200 33 58 109 160 45 107 55 MetaMask (Met, 2023b, a) 200 115 49 40 159 91 70 30 Firefox Nightly (Fir, 2023d, c) 200 38 84 82 6,077 5,243 854 346 Firefox Focus (Fir, 2023b, a) 200 37 78 87 735 597 139 44 SUM 1000 313 332 373 10,217 8,536 1,697 768
We employed the TFAutoModelForSequenceClassification class (TFA, 2023) from Huggingface’s Transformers library to fine-tune pre-trained models. We trained the classification models for 3 epochs using a batch size of 16 and a learning rate of 5e-5 for the Adam optimizer (Kingma and Ba, 2014). Furthermore, we set the maximum input string length to 50, consistent with our datasets. All models were executed on Google Colab, which offers a free GPU with 12.7 GB of system RAM, 15.0 GB of GPU RAM, and 78.2 GB of disk storage. Also, our investigation, which corresponds to RQ4, demonstrates that the DistilBERT model achieves the highest F1-score in the average of the bug report and feature request results with the shortest time for training and prediction phases. So, we used it to train different review classifiers in preceding experiments.
4.3. Evaluation Measures